OpenStudy (anonymous):

What is 211.7647 as a fraction?

OpenStudy (zepp):

What calculator do you have?

OpenStudy (anonymous):

A normal one and a scientific

OpenStudy (zepp):

Um, great, apparently, this number isn't something cool: http://www.wolframalpha.com/input/?i=211.7647

OpenStudy (lgbasallote):

here's a hint: \[50.28 \implies 50 \frac{28}{100} \implies 50 \frac{28 \div 4}{100 \div 4} \implies 50 \frac{7}{25}\] does that help?

OpenStudy (lgbasallote):

or you can use scientific calculator and do 211.7647 then press equals then press that thingy that looks like a b/c

OpenStudy (anonymous):

Ah, yes. Just give me sec to figure it out.

OpenStudy (anonymous):

Thank you guys :).

OpenStudy (lgbasallote):

you're welcome :D